You are here: Foswiki>Commissioning Web>Logs>09Apr2012_Log (09 Apr 2012, JuanGuerra)Edit Attach
-- JuanGuerra - 09 Apr 2012

UT 20120409 AO#2 Commissioning

Observer: ARiccardi, LBusoni, FQuiros-Pacheco, CArcidiacono (LBTO)
Telescope Operator: GBechetti (LBTO)
Additional AO Support: JCGuerra
Telescope Support: JLittle(LBTO)
Instrument Support: DOfficer (LBTO)

Summary

Cirrus passing.

Close dome tests. Retro-reflector on to identify the problem of the last night.

Start up of the systems. W# OK, Secondary problems during the power on.

Secondary right and W#1 power on and set. Not problems.

Many problems with the secondary showing a low stability. At some point we lost the signal of one of the crates (see screenshot)

Details

14:00 Retro- installed in the telescope and systems are started up.

System power on and checking a bin1 without disturbance, the system seems to work, BUT, the secondary start to RIP during the close loop and during the recover failure.

Starting to dig into the logs to identify the cause of the continous failures.

Tescope at horizon to open the secondary cabinets (could be a temperature issue?). This is not a direct effect to the failures and the temperatures were normal before to open the cabinet.

Secondary is set without the FF (apply a simple flat) and the system is tested at close loop after two minutes the shell RIP. The shell is recovered and it was RIP during the process of set flat.

The shell RIP even in set position (many times).

2:52 We are still trying to identify the cause of the RIPs .

It was started working with the Bin1 to determine if we have problems with the reconstructors. The shell started to RIP even during the set flat or in rest position and at some point one of the crates were off. analysing the reason why this sector of actuators jump (same bcu from last replacement), we decided to go up to the secondary and try to explore the possibility to change the signal generator board. Looking into the secondary and trying to remove the board it has seen that the board ca not be removed because it is too inside in the hub and in the case to remove this board we need to do again the optical calibration of the secondary.

08:00 The secondary system is restarted. The secondary seems more stable than before and it gives the opportunity to work with different close loop and take some diagnostic data.

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120409_082130 ??? 7.3 90 6 NaN -1.00 180658 1 400 990 0.4 0.2 0.0 3 693 ON 60.4 J 20 200 OFF NO  
20120409_082158 ??? 7.3 90 8 NaN -1.00 180658 1 400 990 0.4 0.2 0.0 3 690 ON 59.0 J 20 200 OFF NO  
20120409_082221 ??? 7.3 90 8 NaN -1.00 180658 1 400 990 0.4 0.2 0.0 3 687 ON 62.0 J 20 200 OFF NO  
20120409_082627 ??? 7.4 90 8 NaN -1.00 180658 1 400 990 0.4 0.2 0.0 3 679 ON 53.8 J 20 200 OFF NO  
20120409_082648 ??? 7.3 90 8 NaN -1.00 180658 1 400 990 0.4 0.2 0.0 3 684 ON 55.8 J 20 200 OFF NO  
20120409_084633 ??? 7.3 90 4 NaN -1.00 181326 1 500 990 0.5 0.2 0.0 3 687 ON 60.7 J 20 200 OFF NO  
20120409_084653 ??? 7.4 90 6 NaN -1.00 181326 1 500 990 0.5 0.2 0.0 3 680 ON 62.3 J 20 200 OFF NO  
20120409_084714 ??? 7.4 90 6 NaN -1.00 181326 1 500 990 0.5 0.2 0.0 3 680 ON 61.5 J 20 200 OFF NO  
20120409_084902 ??? 7.4 90 3 NaN -1.00 181326 1 500 990 0.5 0.2 0.0 3 680 ON 61.8 J 20 200 OFF NO  
20120409_084921 ??? 7.4 90 5 NaN -1.00 181326 1 500 990 0.5 0.2 0.0 3 681 ON 59.0 J 20 200 OFF NO  
20120409_084940 ??? 7.3 90 6 NaN -1.00 181326 1 500 990 0.5 0.2 0.0 3 688 ON 58.9 J 20 200 OFF NO  
20120409_085027 ??? 7.4 90 7 NaN -1.00 181326 1 500 990 0.5 0.2 0.0 3 675 ON 60.1 J 20 200 OFF NO  
20120409_085045 ??? 7.4 90 5 NaN -1.00 181326 1 500 990 0.5 0.2 0.0 3 683 ON 62.4 J 20 200 OFF NO  
20120409_085104 ??? 7.4 90 8 NaN -1.00 181326 1 500 990 0.5 0.2 0.0 3 683 ON 59.1 J 20 200 OFF NO  
20120409_085532 ??? -NaN 90 3 NaN -1.00 181326 1 500 990 0.5 0.2 0.0 3 -1 ON 60.7 J 20 200 OFF NO  
20120409_085550 ??? -NaN 90 2 NaN -1.00 181326 1 500 990 0.5 0.2 0.0 3 -1 ON -1.0 ? -1 -1 OFF NO  

The loop was closed with the 400 modes and we could take some optical loop data. There is a difference of the performances from the last time that the system run the 400 modes (three days ago ~73%, today ~60%).

Test with 500 modes and it has seen that the loop is open during the gain optimization. Different test has been done to understand why the system is instable for 500 modes and the secondary started to be instable (giving many RIPs). Recover the secondary and we went back to close the loop (400 modes)

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120409_111439 ??? 7.4 90 3 NaN -1.00 180658 1 400 990 0.3 0.5 0.3 3 679 ON 44.7 J 20 200 ON NO  
20120409_111458 ??? 7.3 90 4 NaN -1.00 180658 1 400 990 0.3 0.5 0.3 3 688 ON 46.8 J 20 200 ON NO - SR may be in error
20120409_111516 ??? 7.3 90 2 NaN -1.00 180658 1 400 990 0.3 0.5 0.3 3 688 ON 46.1 J 20 200 ON NO  
20120409_111609 ??? 7.4 90 2 NaN -1.00 180658 1 400 990 0.3 0.5 0.3 3 650 ON 44.4 J 20 200 ON NO  
20120409_111627 ??? 7.3 90 5 NaN -1.00 180658 1 400 990 0.3 0.5 0.3 3 685 ON 45.6 J 20 200 ON NO  
20120409_111645 ??? 7.3 90 4 NaN -1.00 180658 1 400 990 0.3 0.5 0.3 3 687 ON 48.8 J 20 200 ON NO - SR may be in error
20120409_111934 ??? 7.4 90 4 NaN -1.00 180658 1 400 990 0.3 0.5 0.3 3 683 ON 45.2 J 50 50 ON NO  
20120409_111945 ??? 7.3 90 5 NaN -1.00 180658 1 400 990 0.3 0.5 0.3 3 694 ON 46.2 J 50 50 ON NO  
20120409_111957 ??? 7.3 90 6 NaN -1.00 180658 1 400 990 0.3 0.5 0.3 3 691 ON 46.8 J 50 50 ON NO  
20120409_112032 ??? 7.4 90 2 NaN -1.00 180658 1 400 990 0.3 0.5 0.3 3 651 ON 45.5 J 50 50 ON NO  
20120409_112044 ??? 7.4 90 2 NaN -1.00 180658 1 400 990 0.3 0.5 0.3 3 681 ON 47.0 J 50 50 ON NO  
20120409_112056 ??? 7.3 90 6 NaN -1.00 180658 1 400 990 0.3 0.5 0.3 3 684 ON 48.3 J 50 50 ON NO  

After a few more RIPs of the shell without any aparent reason and suspecting that there is something wrong in the hardware we decided to stop.

Stop 12:32

  • crate-failure:
    crate-failure2.png
Topic revision: r4 - 09 Apr 2012, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback